# Advancing Large Multi-modal Models with Explicit Chain-of-Reasoning and Visual Question Generation

### v1

@@ 4.3 Stage3: CoR Fine-tuning
When the model is applied to test data, it performs a two-stage reasoning process to reach the final answer. The first stage encompasses the reasoning that leads up to the generation of a question. The second stage occurs after an external response to the question is obtained; the model then generates the remaining reasoning steps based on this new information. In this study, we leverage GPT-4 as the external answerer. It is noteworthy that the questions generated by the model typically demand an understanding of the image content as well as common sense knowledge. Given its extensive training data and large model parameters, GPT-4 is expected to adeptly handle the common sense aspect. For image comprehension, just as during dataset creation, we furnish GPT-4 with image captions, object classes, and region coordinates, thus effectively providing it with the necessary image context to contribute informed responses.
@@ 5.1 Implementation Details
Our experimental setup incorporates the following models: for image processing, we use CLIP-ViT-Large, and for text generation, we employ LLaMA2-chat-13b.
In stage 1 training, we utilized 4 Tesla A100 (80GB) GPU for 10.5 days. We trained the model for one epoch with a total batch size of 512. The learning rate was set to a maximum of 1e-5, linearly increasing until 20% of the training steps, followed by a cosine decay schedule. Training in Stage 2 was conducted for about 8 days using 8 A100 GPUs (80GB). The training consisted of 30,000 steps. We set the total batch size to 192, the maximum learning rate to 5e-5, increasing linearly for the first 5,000 steps with same cosine decay schedule. In the final stage, we utilized 8 A100 (80GB) GPU. The model was run for 10 epochs, and it took about 4 hours to complete. The batch size remained consistent at a total of 192. Same as before steps, the learning rate was linearly increased to 1e-5 until 20% of the training steps, followed by a cosine decay schedule. Entire training was conducted using bfloat16 precision.
@@ 5.2 Evaluation Settings
In our experiments, we investigated how the presence or absence of CoR affects the behavior of the model when solving tasks. We used validation set of four datasets for this evaluation: MS COCO [22], Visual Genome [20], OK-VQA [31], and A-OKVQA [42]. Our experiments compared the model trained with CoR to a baseline model trained without question generation, specifically using the “with GT” setting from the dataset (referred to as “w/o Question”). To extract the final answer from the generated CoR, we employed GPT-4, guided by the prompt: Extract the final answer from the below reasoning step. The response must only include the answer, in a single word or a short phrase. If you cannot extract the answer, please write ‘No answer’. Additionally, we also evaluated the outputs generated by a model that did not undergo stage 3 training, meaning it was not trained on the CoR dataset. These results were labeled as “w/o CoR” in the results.
For evaluation metrics, we applied BLEU [37], METEOR [4], and CIDEr [50] for captioning tasks (MS COCO, Visual Genome), and the VQA score [3] for VQA tasks (OK-VQA, A-OKVQA). We used nlg-eval library111https://github.com/Maluuba/nlg-eval for the calculation of BLEU, METEOR, and CIDEr.
@@ 5.3 Results and Discussions
The results of our experiments are shown in Table 2. Also, we show an example of the predicted reasoning steps in Figure 5. First, we compare the performance of models utilizing CoR. Comparing the “w/o Question” and “with Question” settings, we observed that the latter yielded better results in most metrics, except for the METEOR score in Visual Genome and the accuracy in A-OKVQA.
The improvement with the “with Question” setting underscores the importance of generating questions and acquiring information during the reasoning process. Notably, the performance gain in OK-VQA was significant, at approximately 32%. Given that OK-VQA includes questions that often require knowledge beyond the content of the image, such as common-sense, it suggests that the process of acquiring external knowledge through asking questions effectively complemented the image-based reasoning steps. However, when compared to models not using CoR, we noticed a decrease in evaluation scores for models using CoR. This indicates that current models may struggle to accurately generate long reasoning steps and consistently reach the final answer. The challenge of developing LMMs that can produce coherent and consistent long reasoning steps in line with the given tasks remains an important area for future research.
@@ 6 Conclusion
In this paper, we have presented a novel approach of improving the capabilities of LMMs. This was achieved by incorporating a structured CoR and the ability to generate questions during the reasoning process. The model was trained on a dataset specifically designed to include an explicit reasoning and question-asking process.
We proposed an architecture for an LMM that leverages pre-trained components such as an image encoder and text decoder, all of which are fine-tuned using our novel dataset. Our experimental results suggest that the model’s ability to generate questions contributes to its performance, compared to a baseline model that does not generate questions. This underscores the potential utility of integrating explicit reasoning processes into LMMs.
@@ Acknowledgements
More Examples of the Dataset
More Qualitative Results
We show more qualitative results of generated reasoning steps by our model for MS COCO, Visual Genome caption, OK-VQA, and A-OKVQA in Figure 8 and 9.

### v2

@@ 3.4 Training
When the model is applied to test data, it performs a two-stage reasoning process to reach the final answer. The first stage encompasses the reasoning that leads up to the generation of a question. The second stage occurs after an external response to the question is obtained; the model is given the original inputs, the intermediate reasoning steps, and the external response. Then, the model generates the remaining reasoning steps based on this new information.
In this study, we leverage GPT-4o as the external answerer. It is noteworthy that the questions generated by the model typically demand an understanding of the image content as well as common-sense and encyclopedic knowledge. Given its extensive training data and large model parameters, GPT-4 is expected to adeptly handle the knowledge required to answer the questions.
@@ 4.1 Implementation Details
Our experimental setup incorporates the following models: for image processing, we use CLIP-ViT-Large111https://huggingface.co/openai/clip-vit-large-patch14-336, and for text generation, we employ vicuna-7b-1.5222https://huggingface.co/lmsys/vicuna-7b-v1.5. The entire model is trained from a publicly available pre-trained checkpoint 333https://huggingface.co/llava-hf/llava-1.5-7b-hf.
In fine-tuning stage, we utilized 8 NVIDIA H100 (80GB) GPU for about 1 hour. We trained the model for one epoch with a total batch size of 128. The learning rate was set to a maximum of 2e-5, linearly increasing until 3% of the training steps, followed by a cosine decay schedule. Training was conducted using bfloat16 precision.
@@ 4.2 Evaluation Settings
In our experiments, we investigated how the presence or absence of CoR affects the behavior of the model when solving tasks. We also evaluated several ablations related to the CoR method. These included listing the necessary knowledge before question generation and generating an uncertainty score along with the reasoning steps. We used validation set of seven datasets for this evaluation: MS COCO, Visual Genome, VQAv2, OK-VQA, A-OKVQA, OVEN, and Encyclopedic VQA. Using entire validation set is computationally expensive, so we randomly sampled 1000 instances from each dataset for evaluation. Our experiments compared the model trained with CoR to a baseline model trained without question generation, specifically using the “with GT” setting from the dataset (referred to as “w/o Question”).
Currently, there are no established metrics for evaluating results that include long reasoning steps. Therefore, in this study, we used GPT-4 for scoring the answers. We instructed GPT-4 to extract the final answer from the reasoning steps and then evaluate the answer with a score from 1 to 4 (1: incorrect, 2: partially correct, 3: mostly correct, 4: correct).
@@ 4.3 Results and Discussions
The results of our experiments are shown in Table 2. Also, we show an example of the predicted reasoning steps in Figure 6.
The evaluation results show that “Ours CoR” performs the best on average across various datasets. In contrast, the CoR model without questions (CoR w/o question) has the lowest performance. This might be because the current model cannot consistently provide coherent outputs over long reasoning processes. The challenge of developing VLMs that can produce coherent and consistent long reasoning steps in line with the given tasks remains an important area for future research.
Comparing “Ours CoR” with the original LLaVA, the latter performs relatively well on VQA v2 and COCO Caption, which are part of LLaVA’s training data.However, for datasets not included in LLaVA’s training data, like OVEN and Encyclopedic VQA, which require specialized knowledge, models that ask questions (such as “Ours CoR”) show significantly better performance. This indicates that asking questions helps the model acquire necessary specialized knowledge that it does not inherently possess.
Furthermore, outputting uncertainty did not significantly improve performance (CoR w/ uncertainty vs. Ours CoR). This is likely because generating uncertainty outputs makes the model’s responses longer and less consistent, leading to performance degradation. However, if the model can handle uncertainty effectively, it could ask questions during the most uncertain reasoning steps, potentially improving performance.
Comparing models with and without prior knowledge output (CoR w/o knowledge vs. Ours CoR), it’s evident that pre-outputting necessary knowledge improves performance. This prevents the model from asking irrelevant questions and ensures that it focuses on helpful inquiries, leading to better results.
@@ 5 Conclusion
In this paper, we have presented a novel approach of improving the capabilities of VLMs. This was achieved by incorporating a structured CoR and the ability to generate questions during the reasoning process. The model was trained on a dataset specifically designed to include an explicit reasoning and question-asking process.
We proposed an architecture for a VLM that leverages pre-trained components such as an image encoder and text decoder, all of which are fine-tuned using our novel dataset. Our experimental results suggest that the model’s ability to generate questions contributes to its performance, compared to a baseline model that does not generate questions. This underscores the potential utility of integrating explicit reasoning processes into VLMs.
@@ Acknowledgements
A. Prompt Details
Here, we provide the full example of the prompt given to GPT-4 for answer generation and evaluation. The prompt for answer generation is shown in Listing LABEL:lst:prompt_answer, and the prompt for answer evaluation is shown in Listing LABEL:lst:prompt_eval. In both prompts, we provide an instruction and several examples to guide GPT-4 to generate and evaluate the answer, respectively. We instructed GPT-4 to return the answer in JSON format, which enables us to parse the answer and evaluate it automatically.
B. More Examples of the Dataset
In Figure 7 and 8, we show more examples of the dataset we crafted based on different datasets: OK-VQA, A-OKVQA, MS COCO caption, Visual Genome caption, OVEN, and Encyclopedic-VQA. Each example consists of an image, a question, and a reasoning chain that explains how to answer the question or generate a caption for the image. We provide the reasoning chain of three types: (1) reasoning steps without answer and question, in which the data is generated without access to the ground-truth answer and there is no question generation step; (2) reasoning steps with question, in which the data is generated with access to the ground-truth answer and there is a question generation step; and (3) reasoning steps with answer and without question, same as the second type but without the question generation step.
C. More Qualitative Results
We show more qualitative results of generated reasoning steps by our model for VQA, A-OKVQA, COCO Caption, Encyclopedic VQA, OVEN, OK-VQA, and Visual Genome in Figure 10, 11, 12, and 13.